# GAVEL: Generating Games Via Evolution and Language Models

### v1

@@ 1 Introduction
Games have long been used as a test bed for algorithms and approaches in artificial intelligence, with advances in game-playing ability often serving as some of the most recognizable achievements in the field [12, 39, 53, 59]. While automated systems have repeatedly demonstrated an ability to match or surpass humans as game players, they continue to lag significantly in their capacity to generate the kinds of games that are worth playing. The ability to construct novel and interesting games is an impressive cognitive challenge, and success would have implications both cultural (with the production of new artifacts) and computational (with the generation of new learning environments for artificial agents). Prior efforts in automated game design [41, 57] have produced some successes (notably the commercially-available game Yavalath, generated by the Ludi system [11]), but remain largely limited by hard-coded heuristics and restricted domains. These limitations are often necessary, however, in the face of automated game design’s core challenges: representing the vast array of possible games in a structured and computationally workable form, and efficiently searching through the resulting representation space for worthwhile games.
@@
We show empirically that GAVEL is capable of generating playable and interesting board games that differ substantially from games encountered during training. Our approach intelligently recombines mechanics and ideas from disparate genres and produces samples that mirror the performance of human-generated games under a suite of automated evaluation metrics. A preliminary qualitative analysis also reveals that GAVEL can generate novel games that are both engaging and entertaining. We conclude with a discussion of GAVEL’s successes and failures, as well as the promising avenues for future work.
@@ 2.1 Automatic Game Generation
Our work continues a strand of research that investigates the ability for automated systems to produce novel games or novel variants of existing games. The first such effort was METAGAME [44], which generates games by sampling from a grammar that encodes “symmetric Chess-like games” with designer-specified rule probabilities. Since then, work has continued in the generation of both board game and video game rulesets. Evolutionary or search-based game design is a popular technique, as game descriptions do not typically afford gradient information; it was first proposed in 2008 for board games [11] and video games [57], and later work has brought it to bear on different video game genres [42, 15, 28]. Another approach begins instead from conceptual or symbolic specifications of rules or mechanics and generates games by dynamically referring to a pre-specified library of gameplay elements [41, 58]. Yet another approach is to use constraint satisfaction algorithms. For instance, rules might be encoded as an answer-set program, with constraints pre-specified by a designer to define what counts as an acceptable game [54, 64, 43, 25, 56]. Most recently, work has investigated the ability for large language models to act as design assistants by proposing game mechanics [2] or directly synthesizing small programs [26].
@@ 2.2 Evolutionary Computation and Language Models
In recent years, improvements in large language models both generally [6, 63] and in their ability to produce code [13, 61] have led to their use in a variety of evolutionary systems. Of note is evolution through large models [33], which learns a diff model (i.e. a language model that can modify programs conditioned on a natural language specification) from a dataset of GitHub commits and accompanying messages. This model is then used as the mutation operator for genetic programming in Python through a quality-diversity algorithm [40]. We adopt this general approach, though GAVEL learns to mutate games from raw programs (i.e. without diffs or commit messages) in a domain-specific language. Similar techniques have also been used to generate programming puzzles [47], reward functions for reinforcement learning [37], and poetry [5].
@@ 3.1 Dataset
After this, we filter our dataset to remove a small number of puzzles, experimental games and Mancala-style games, as their structure and general gameplay rules are very dissimilar from the rest of the dataset. We then tokenize each game according to our code language model (see below) and exclude any game that is longer than 1024 tokens. From this reduced dataset, we hold out a set of 14 varied games (available in Appendix A) that will be used to initialize the evolutionary search, with the remaining 574 games being used as our training dataset. An example of one of the held-out games and its converted form is available in Figure 2.
@@ 4.1 Language Model Training
We note here that FITM training may have a potential downside in the context of evolution through large models. Specifically, the model is trained to perfectly reproduce the missing section of code given its prefix and suffix. If it succeeds completely in doing so during evolution, then the resulting game will not be mutated at all. In essence, there is a tension between the ability of the model to accurately capture the underlying logic and syntax of the representation space and its ability to memorize or perfectly reconstruct its training dataset. In GAVEL, we mostly sidestep this issue by mutating a set of held-out games not seen at all during training (making memorization impossible), though see Section 8 for a discussion of other possible approaches.
@@ 4.2 Evolutionary Search
Determining an appropriate set of behavioral characteristics is challenging in the context of automatic game generation. Intuitively, distinct archive cells ought to capture meaningfully distinct games while collapsing minor or trivial variations. Human game players readily make such categorizations, but automatically identifying such differences (especially with previously unseen games) is both difficult and necessary for the MAP-Elites algorithm to function. In order to tackle this problem, we take advantage of the semantic concepts described in Section 3. Building on evidence that Ludii concept vectors capture a meaningful notion of distance [55], we use principal component analysis (PCA) [60] on the complete Ludii dataset (i.e. before any filtering) to reduce the 510-dimensional concept vectors to two dimensions. We then bucket the resulting two-dimensional space into 40 equally-spaced regions from -5 to 5 in each dimension, obtaining a rectangular archive of 1600 cells. While the first two PCA dimensions describe only of the variance in the concept feature space, a preliminary investigation indicated that increasing the number of dimensions resulted in a less diverse archive. Because games are not uniformly distributed through feature space, increasing the archive’s dimensionality for a fixed number of cells caused a larger number of distinct games to be mapped to the same cell. See discussion of potential alternatives in Section 7.
We initialize the archive by adding and evaluating the 14 heldout games listed in Appendix A. For each MAP-Elites step, we select games from the current archive. For each game, we then select random parenthetical expressions and re-format them as a (prefix, suffix, target) tuple. We then sample from the trained CodeLlama-13b model with a temperature of 1 to generate a new expression, conditioned on just the prefix and suffix, and re-construct the resulting game. After filtering out any duplicate or unchanged games, the samples are evaluated for fitness and assigned an archive cell based on the PCA reduction of their concept vector.
@@ 6.1 Quantitative Results
All Cells | Novel Cells | ||||
Run | QD Score | # Playable | # Fitness>0.5 | # Playable | # Fitness>0.5 |
Overall, GAVEL succeeds in generating a wide range of novel and high-fitness games. In Table 1 we present the mean and standard deviations of the archive metrics for GAVEL and GAVEL-UCB. The quality diversity score is difficult to interpret in isolation, as its magnitude depends greatly on the potential range of fitness scores. In this case, it is most helpful as a way to compare the performance of disparate algorithms on the same task: we see that GAVEL improves significantly over GAVEL-UCB (Welch’s -test, ), indicating that it has some mixture of higher fitness and greater variety in the samples it produces. Of more interest is the fact that both methods manage to fill a substantial proportion of the archive with playable games while only starting from a modest 14 samples, including regions of the archive not covered by games in the Ludii training dataset (differences between GAVEL and GAVEL-UCB are not significant, ). We present a visualization of the archive produced by one run of GAVEL over time in Figure 3, which shows both the success of the model in generating high-fitness samples and the fact that much of the concept space remains unexplored.
@@ 6.2 Qualitative Results
Figure 2 (right) includes another example of a generated game noted by our evaluators to be particularly interesting. It is a variant of Havannah (a game in which players attempt to form loops or connected lines of pieces between sides of the board) that introduces a restriction on piece placement from another game in the Ludii dataset (Tabu Y). A final exemplar, presented in Appendix C alongside the previous two examples, modifies the pawn-advancement game Breakthrough to use pieces that can only move by hopping over each other (and without capturing). Taken together, these examples demonstrate the strength of GAVEL: it is able to intelligently recombine game mechanics (expressed as code segments) in novel ways and ensure that these combinations do not result in trivial games.
@@ 7 Discussion and Limitations
Heldout Games: as noted in Section 4.2, we initialize the search using a set of games held out from the language model training in order to prevent memorization. While GAVEL produces a wide range of games from this set, they nonetheless will share many features with one another as a result of their shared origins. One potential solution to further increase archive diversity is modifying the mutation process directly, either by increasing sampling temperature or by enforcing novelty with respect to the tokens in the original game section through masking. These techniques might also make it possible to initialize the archive with games in the training dataset, further increasing potential output diversity.
@@ Appendix B Language Model Training Hyperparameters
Appendix C GAVEL Exemplars

### v2

@@ 1 Introduction
Games have long been used as a test bed for algorithms and approaches in artificial intelligence, with advances in game-playing ability often serving as some of the most recognizable achievements in the field [13, 40, 55, 65]. While automated systems have repeatedly demonstrated an ability to match or surpass humans as game players, they continue to lag significantly in their capacity to generate the kinds of games that are worth playing. The ability to construct novel and interesting games is an impressive cognitive challenge, and success would have implications both cultural (with the production of new artifacts) and computational (with the generation of new learning environments for artificial agents). Prior efforts in automated game design [42, 62] have produced some successes (notably the commercially-available game Yavalath, generated by the Ludi system [12]), but remain largely limited by hard-coded heuristics and restricted domains. These limitations are often necessary, however, in the face of automated game design’s core challenges: (1) representing the vast array of possible games in a structured and computationally workable form and (2) efficiently searching through the resulting representation space for worthwhile games.
@@
We show empirically that GAVEL is capable of generating playable and interesting board games that differ substantially from games encountered during training. Our approach intelligently recombines mechanics and ideas from disparate genres and produces samples that mirror the performance of human-generated games under a suite of automated evaluation metrics. A preliminary qualitative analysis also reveals that GAVEL can generate novel games that are both engaging and entertaining. We conclude with a discussion of GAVEL’s successes and failures, as well as the promising avenues for future work. We provide a public repository that includes our code and data, including a trained model checkpoint. 222Code and data available here: https://github.com/gdrtodd/gavel
@@ 2.1 Automatic Game Generation
Our work continues a strand of research that investigates the ability for automated systems to produce novel games or novel variants of existing games. The first such effort was METAGAME [45], which samples from a grammar that encodes “symmetric Chess-like games” with designer-specified rule probabilities. Since then, work has continued in the generation of both board game and video game rulesets. Evolutionary or search-based game design is a popular technique, as game descriptions do not typically afford gradient information; it was first proposed in 2008 for board games [12] and video games [62], and later work has brought it to bear on different video game genres [43, 16, 29]. Another approach begins instead from conceptual or symbolic specifications of rules or mechanics and generates games by dynamically referring to a pre-specified library of gameplay elements [42, 63]. Yet another approach is to use constraint satisfaction algorithms. For instance, rules might be encoded as an answer-set program, with constraints pre-specified by a designer to define what counts as an acceptable game [56, 71, 44, 26, 59]. Most recently, work has investigated the ability for large language models to act as design assistants by generating game levels [61, 58], proposing game mechanics [3] or directly synthesizing small programs [27].
@@ 2.2 Evolutionary Computation and Language Models
In recent years, improvements in large language models both generally [7, 70] and in their ability to produce code [14, 67] have led to their use in a variety of evolutionary systems. Of note is evolution through large models [34], which learns a diff model (i.e. a language model that can modify programs conditioned on a natural language specification) from a dataset of GitHub commits and accompanying messages. This model is then used as the mutation operator for genetic programming in Python through a quality-diversity algorithm [41]. We adopt this general approach, though GAVEL learns to mutate games from raw programs (i.e. without diffs or commit messages) in a domain-specific language. Similar techniques have also been used to generate reinforcement learning environments [1, 69] and reward functions [38], adversarial prompts [53], programming puzzles [48], and poetry [6].
@@ 3.1 Dataset
After this, we filter our dataset to remove a small number of puzzles and experimental games. We also remove Mancala-style games, as prior work has identified them as occupying a very distinct cluster with respect to the full collection of Ludii games, in terms of rules and structure [57]. We then tokenize each game according to our code language model (see below) and exclude any game that is longer than 1024 tokens. From this reduced dataset, we hold out a set of 14 varied games (available in Appendix A) that are used to initialize the evolutionary search, with the remaining 574 games being used as our training dataset. An example of one of the held-out games and its converted form is available in Figure 2.
@@ 4.1 Language Model Training
We note here that FITM training may have a potential downside in the context of evolution through large models. Specifically, the model is trained to perfectly reproduce the missing section of code given its prefix and suffix. If it succeeds completely in doing so during evolution, then the resulting game will not be mutated at all. In essence, there is a tension between the ability of the model to accurately capture the underlying logic and syntax of the representation space and its ability to memorize or perfectly reconstruct its training dataset. In GAVEL, we mostly sidestep this issue by mutating a set of held-out games not seen at all during training (making memorization impossible), though see Appendix C for an initial investigation on mutating training-set games and Section 8 for a discussion of other possible approaches.
@@ 4.2 Evolutionary Search
Determining an appropriate set of behavioral characteristics is challenging in the context of automatic game generation. Intuitively, distinct archive cells ought to capture meaningfully distinct games while collapsing minor or trivial variations. Human game players readily make such categorizations, but automatically identifying such differences (especially with previously unseen games) is both difficult and necessary for the MAP-Elites algorithm to function. In order to tackle this problem, we take advantage of the semantic concepts described in Section 3. Building on evidence that Ludii concept vectors capture a meaningful notion of distance [57], we use principal component analysis (PCA) [66] on the complete Ludii dataset (i.e. before any filtering) to reduce the 510-dimensional concept vectors to two dimensions. We then bucket the resulting two-dimensional space into 40 equally-spaced regions from -5 to 5 in each dimension, obtaining a rectangular archive of 1600 cells. While the first two PCA dimensions describe only of the variance in the concept feature space, a preliminary investigation indicated that increasing the number of dimensions resulted in a less diverse archive (see Appendix D for additional details). Because games are not uniformly distributed through feature space, increasing the archive’s dimensionality for a fixed number of cells caused a larger number of distinct games to be mapped to the same cell. See discussion of potential alternatives in Section 7.
We initialize the archive by adding and evaluating the 14 heldout games listed in Appendix A. For each MAP-Elites step, we select games from the current archive. For each game, we then select random parenthetical expressions and re-format them as a (prefix, suffix, target) tuple. We then sample from the trained CodeLlama-13b model with a temperature of 1 and a top- value of 50 to generate a new expression, conditioned on just the prefix and suffix, and re-construct the resulting game. After filtering out any duplicate or unchanged games, the samples are evaluated for fitness and assigned an archive cell based on the PCA reduction of their concept vector.
@@ 5 Experiments
All Cells | Novel Cells | ||||
Method | QD Score | # Playable | # Fitness>0.5 | # Playable | # Fitness>0.5 |
Pure Sampling | 296.92 ± 14.84 | 89.00 ± 5.66 | 83.00 ± 5.10 | 14.67 ± 3.30 | 11.33 ± 2.87 |
GPT-4o | 268.16 ± 17.33 | 84.67 ± 6.60 | 80.67 ± 5.19 | 16.67 ± 3.30 | 15.33 ± 2.87 |
@@
We compare GAVEL against two baselines: a pure sampling approach, which uses the same fine-tuned large language model but omits the quality diversity search, and direct sampling from GPT-4o. For the pure sampling baseline, we randomly select a game from the validation set and mutate it by re-generating a random parenthetical expression. We repeat this process 4500 times in order to match the number of samples produced by GAVEL in 500 generations with and . We then evaluate the set of 4500 samples for fitness and determine the cell that each game would occupy based on its concepts, allowing us to construct a simulated archive (i.e. by retaining the highest-fitness sample in each cell) and compare directly against GAVEL. We perform the pure-sampling baseline experiment three times with random seeds . For the GPT-4o baseline, we provide the model with each of the 14 validation games as part of the context window and then ask it to create a modification of one randomly-selected validation game (see Appendix E for prompt). We similarly generate 4500 samples and construct a simulated archive in order to facilitate comparisons with GAVEL.
@@ 6.1 Quantitative Results
Overall, GAVEL succeeds in generating a wide range of novel and high-fitness games. In Table 1 we present the mean and standard deviations of the archive metrics for GAVEL, GAVEL-UCB, and our baseline methods. The quality diversity score is difficult to interpret in isolation, as its magnitude depends greatly on the potential range of fitness scores. In this case, it is most helpful as a way to compare the performance of disparate algorithms on the same task: we see that GAVEL improves significantly over GAVEL-UCB (Welch’s -test, ), indicating that it has some mixture of higher fitness and greater variety in the samples it produces. Similarly, GAVEL improves significantly in terms of QD score over both the pure sampling baseline () and the GPT-4o baseline ().
Of more interest is the fact that GAVEL fills a substantial proportion of the archive with playable games despite starting from a modest 14 samples, including in regions of the archive not covered by games in the Ludii training dataset. Compared to both baselines, GAVEL occupies significantly more cells in the archive and produces more high-fitness (i.e. ) samples ( for pure sampling, for GPT-4o), though the differences between GAVEL and GAVEL-UCB are not significant (). While GAVEL also appears to occupy a larger number of novel cells with playable and high-fitness games compared to the baselines, these improvements are not statistically significant at only three runs. We present a visualization of the archive produced by one run of GAVEL over time in Figure 3, which shows both the success of the model in generating high-fitness samples and the fact that much of the concept space remains unexplored.
@@ 6.2 Qualitative Results
Figure 2 (right) includes another example of a generated game noted by our evaluators to be particularly interesting. It is a variant of Havannah (a game in which players attempt to form loops or connected lines of pieces between sides of the board) that introduces a restriction on piece placement from another game in the Ludii dataset (Tabu Y). A final exemplar, presented in Appendix F alongside the previous two examples, modifies the pawn-advancement game Breakthrough to use pieces that can only move by hopping over each other (and without capturing). Taken together, these examples demonstrate the strength of GAVEL: it is able to intelligently recombine game mechanics (expressed as code segments) in novel ways and ensure that these combinations do not result in trivial games.
@@ 7 Discussion and Limitations
Heldout Games: as noted in Section 4.2, we initialize the search using a set of games held out from the language model training in order to prevent memorization. While GAVEL produces a wide range of games from this set, they nonetheless will share many features with one another as a result of their common origins. One potential solution to further increase archive diversity is to improve the variety of mutations, either by increasing sampling temperature or by enforcing novelty with respect to the tokens in the original game section through masking. These techniques might also make it possible to initialize the archive with games in the training dataset, further increasing potential output diversity.
@@ 9 Broader Impact
Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant DGE-2234660. This research is partially funded by the European Research Council as part of the Digital Ludeme Project (ERC Consolidator Grant #771292). The research collaboration was facilitated by COST Action CA22145 - GameTable, supported by COST (European Cooperation in Science and Technology). The authors would like to thank Rushang Gajjal and Sam Earle for their help and constructive feedback throughout the project.
@@ Appendix B Language Model Training Hyperparameters
Appendix C Effects of Mutating Training Games
We present a preliminary investigation on the feasibility of initializing GAVEL with games from the training dataset rather than the held-out validation set. Using the same fine-tuned CodeLlama-13b language model, we generate 300 mutations by sampling a game from the training set, sampling a parenthetical expression from the game, and generating 3 replacements from the model. We perform the same process for the validation set, producing another 300 mutated games. For each mutation, we then compute whether it is an exact duplicate of the originally-sampled game (i.e. whether it is novel) and whether the resulting game compiles under the Ludii grammar (i.e. whether it is valid). We repeat the entire process for different sampling temperatures and settings of for top- sampling, the results of which are presented in Table 2.
Unsurprisingly, mutations produced from training-set games are much less likely to be novel (since the model may have memorized the exact target sequence during training). The fact that roughly half of validation-set mutations are also duplicates might seem surprising at first, but it is important to note that many sampled parenthetical expressions are very short (e.g. (sites)) and might be the only grammatical possibility at that point the game. Training-set mutations do appear to be compilable more often than validation-set mutations (especially at higher sampling temperatures), but this is likely due at least in part to the increased rates of duplication. The proportion of samples that are both novel and valid (i.e. those that represent potentially useful mutations) tells a clear story: sampling mutations from validation-set games is more efficient than doing so from training-set games. A final point of interest is that increasing sampling temperature appears to improve the efficacy of training-set sampling while decreasing the efficacy of validation-set sampling. This further motivates the possible extensions described in Section 7 and Section 8 – high sampling temperature coupled with grammatical sampling constraints might make training-set mutations a viable way forward.
Training-set Mutations | Validation-set Mutations | ||||||
Temperature | Top-K | Novel | Valid | Novel & Valid | Novel | Valid | Novel & Valid |
0.5 | 20 | 17.33 | 99.33 | 16.67 | 47.33 | 98.67 | 46.00 |
50 | 17.33 | 99.33 | 16.67 | 47.33 | 98.67 | 46.00 | |
1 | 20 | 21.00 | 97.33 | 18.33 | 50.67 | 93.67 | 44.33 |
50 | 21.00 | 97.33 | 18.33 | 49.00 | 94.00 | 43.00 | |
1.5 | 20 | 33.67 | 88.67 | 22.33 | 56.67 | 76.67 | 33.33 |
50 | 35.66 | 85.33 | 21.00 | 54.33 | 73.67 | 28.33 |
Appendix D Effects of Archive Dimensionality and Size
As part of initial experiments for GAVEL, we explored the effect of changing the number of PCA dimensions and the total number of cells used by the MAP-Elites archive on the number of unique cells occupied by training and validation games. In each case, we fit a PCA model with specified number of components on the Ludii dataset using the process described in Section 4.2. For a target number of total archive cells and number of dimensions , we constructed a rectangular archive by separating each axis from -5 to 5 into evenly-spaced regions. Of course, for many choices of and , this process produces an archive with a total size much less than (e.g. for and , the archive would have cells). To combat this, we increased the number of regions in the first dimension as much as possible while keeping the total number of cells no more than (e.g. for and , the number of regions in each dimension is for a total archive size of 1000).
We then simulated adding each of the 574 training games and 14 validation games from our dataset into the archive and measured the number of unique resulting cells. This gives a sense of the “diversity” of the archive – if only a small number of cells are occupied it indicates that the archive fails to distinguish between substantively distinct games. On the other hand, if the number of occupied cells is roughly equal to the number of games, this might indicate that the archive assigns even almost identical games to different cells and could slow the search process. The results of these experiments are presented in Table 3. Our investigation indicated that increasing the dimensionality for fixed target archive size caused the set of training games to be collapsed to a smaller set of cells – a 2-dimensional archive with 2500 cells is roughly as diverse under this formulation as a 3-dimensional archive with 10000 cells. We ultimately decided to use a 2-dimensional archive with roughly 1500 cells because it was the smallest archive that mapped each of the validation games to distinct cells.
Target total archive cells | ||||||||
Input set | Dimension | 100 | 500 | 1000 | 1500 | 2500 | 5000 | 10000 |
Training | 2D | 32 | 105 | 166 | 215 | 268 | 370 | 410 |
3D | 24 | 59 | 74 | 101 | 137 | 191 | 247 | |
4D | 19 | 52 | 58 | 75 | 109 | 128 | 180 | |
5D | 45 | 32 | 48 | 70 | 99 | 98 | 126 | |
Validation | 2D | 9 | 13 | 13 | 14 | 14 | 14 | 14 |
3D | 6 | 8 | 9 | 11 | 11 | 10 | 13 | |
4D | 5 | 7 | 8 | 10 | 9 | 12 | 12 | |
5D | 7 | 7 | 8 | 8 | 9 | 8 | 9 |
We also briefly explored the possibility of using a more sophisticated kind of archive (namely a Centroidal Voronoi Tesselation (CVT) based approach [64]) to combat the high dimensionality of the full concept vector space. The CVT archive is designed to scale to high-dimensional problems by breaking down the search space into a pre-specified number of geometrically homogeneous niches, so we attempted to apply it to the full 510-dimensional Ludii concept vectors using the implementation in the Pyribs library [60] and thereby avoid the need for lossy dimensionality reduction. We set the boundaries in each dimension to because each concept is binary. Unfortunately, even very large archives (i.e. with 100000 cells) collapsed the set of training and validation games into a small number of unique cells. We attribute this to the fact that Ludii games are not uniformly distributed throughout the space of concepts (i.e. many concepts are correlated or mutually exclusive) while the CVT algorithm assigns equal “resolution” to all parts of the search space. However, it is possible that some combination of PCA and the CVT archive might achieve a better balance of archive size and diversity.
Number of archive cells | |||||||||
Input Set | 100 | 500 | 1000 | 1500 | 2500 | 5000 | 10000 | 50000 | 100000 |
Training | 12 | 19 | 34 | 48 | 36 | 37 | 24 | 41 | 68 |
Validation | 3 | 8 | 9 | 7 | 9 | 7 | 4 | 10 | 7 |
Appendix E LLM Baseline Details
As mentioned in Section 5, we use the GPT-4o language model through the OpenAI API. To increase the diversity of generated samples, we use a sampling temperature of 1. We specify a unique seed for each sample from the model (i.e. seeds 0 through 4499 for the first experiment), though the OpenAI API does not guarantee reproducibility for a specific seed. We provide the system prompt and user prompt used in the baseline experiments below.
System Prompt ⬇ You are an expert programming agent in the Ludii game description language. You output syntactically correct Ludii game descriptions and no other text of any kind.
Game Modification Prompt ⬇ Your task is to mutate a game written in the Ludii game description language to produce a new game. Use the following games as reference for proper Ludii syntax and game structure: =====Game 1====== {REFERENCE_GAME_CODE} ========== . . . =====Game i====== {REFERENCE_GAME_CODE} ========== Now, create a modification of the following game. Make sure to obey the constraints of the Ludii grammar to create a syntactically-valid games. In addition, make sure to modify at least part of the game so that it becomes a new game. Do not simply copy an existing game. =====Game to modify===== {GAME_CODE_TO_MODIFY} ========== =====Modified game=====
Appendix F GAVEL Exemplars